OpenStudy (anonymous):

Determine the resultant if a person walks 9 km south then 12 km west.

OpenStudy (anonymous):

Use the Pythagorean theorem. :D Let 9km= a 12 km= b resultant= c \[a^2+b^2= c^2\] Soooo: \[c^2=a^2+ b^2\] \[c= \sqrt{a^2+ b^2}\] Substitute: \[Resultant= \sqrt{9^2 + 12^2}\] \[Resultant= \sqrt{81+144}\] \[Resultant= \sqrt{225}\] Resultant = 15 km

OpenStudy (anonymous):

Put numbers in here: http://hyperphysics.phy-astr.gsu.edu/hbase/vect.html#vec5

OpenStudy (anonymous):

thanks ^^ practically, this is a very easy question :))

OpenStudy (anonymous):

Duuuuh~~ :P